Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Friends Probability Question

Expert replies
by newgmattest » Tue May 31, 2011 4:21 am
In a room with 7 people, 4 people have exactly 1 friend in the room and 3 people have exactly 2 friends in the room (Assuming that friendship is a mutual relationship, i.e. if Jane is Paul's friend, Paul is Jane's friend). If two individuals are selected from the room at random, what is the probability that those two individuals are NOT friends?

1)5/21
2)3/7
3)4/7
4)5/7
5)16/21

OA : E
Join the discussion
Source: — Problem Solving |

by Frankenstein » Tue May 31, 2011 4:50 am
Hi,
Let the friends be 1,2,3,4,5,6,7. According to the problem, 1,2 are friends, 3,4 are friends and 5,6,7 are friends.
Number of ways of picking 2 so that they a re friends is picking(1,2)+picking(3,4)+picking 2out of (5,6,7) = 1+1+3C2 = 5
Total number of ways of picking 2 guys from the group of 7 = 7C2 = 21.
So, probability that the friends picked are friends is 5/21.
Thus, the probability that those two individuals are NOT friends is 1-5/21 =16/21

Hence, answer E

Cheers!
Join the discussion

by Anurag@Gurome » Tue May 31, 2011 5:05 am
newgmattest wrote:In a room with 7 people, 4 people have exactly 1 friend in the room and 3 people have exactly 2 friends in the room (Assuming that friendship is a mutual relationship, i.e. if Jane is Paul's friend, Paul is Jane's friend). If two individuals are selected from the room at random, what is the probability that those two individuals are NOT friends?

1)5/21
2)3/7
3)4/7
4)5/7
5)16/21

OA : E
Probability of choosing 2 people who are not friends out of the 4 with one friend = 4/7 * 5/6
Probability of choosing 2 people who are not friends out of the 3 with two friends = 3/7 * 4/6
Therefore, required probability = 4/7 * 5/6 + 3/7 * 4/6 = 32/42 = 16/21

The correct answer is E.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by newgmattest » Tue May 31, 2011 5:13 am
Hi Anurag. Thanks for reply. Can you please explain how we derived second multiple in each pair?
Join the discussion

by Anurag@Gurome » Tue May 31, 2011 6:47 am
newgmattest wrote:Hi Anurag. Thanks for reply. Can you please explain how we derived second multiple in each pair?
Let us assume that A, B, C, D, E, F, G are 7 people.
Now possible combinations so that 4 have exactly 1 friend and 3 have exactly 2 friends.

A-B
B-A
C-D
D-C
E-F-G
F-E-G
G-E-F

If we choose one of the 4 people with one other friend, then we have 5/6 chance of not picking their 2nd friend.
Similarly, if we choose one of the 3 people with 2 friends, we have 4/6 chance of not picking one of their 2 friends.

Hope that helps.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion